### v2

# Isogeny relations in products of families of elliptic curves

Abstract.
Let $E_{\lambda}$ be the Legendre family of elliptic curves with equation $Y^{2}=X(X-1)(X-\lambda)$. Given a curve $\mathcal{C}$, satisfying a condition on the degrees of some of its coordinates and parametrizing $m$ points $P_{1},\ldots,P_{m}\in E_{\lambda}$ and $n$ points $Q_{1},\ldots,Q_{n}\in E_{\mu}$ and assuming that those points are generically linearly independent over the generic endomorphism ring, we prove that there are at most finitely many points $\mathbf{c}_{0}$ on $\mathcal{C}$, such that there exists an isogeny $\phi:E_{\mu(\mathbf{c}_{0})}\rightarrow E_{\lambda(\mathbf{c}_{0})}$ and the $m+n$ points $P_{1}(\mathbf{c}_{0}),\ldots,P_{m}(\mathbf{c}_{0}),\phi(Q_{1}(\mathbf{c}_{0})),\ldots,\phi(Q_{n}(\mathbf{c}_{0}))\in E_{\lambda(\mathbf{c}_{0})}$ are linearly dependent over $\operatorname{End}(E_{\lambda(\mathbf{c}_{0})})$.
1. Introduction
Let $m,n$ be positive integers and let $E_{\lambda}$ denote the elliptic curve with Legendre equation
$Y^{2}Z=X(X-Z)(X-\lambda Z)$
and consider this as a family of elliptic curves $E_{\lambda}\rightarrow Y(2)=\mathbb{A}^{1}\setminus\left\{0,1\right\}$. With a slight abuse of notation, we will denote by $E_{\lambda}^{m}$ the $m$-fold fibered power $E_{\lambda}\times_{Y(2)}\ldots\times_{Y(2)}E_{\lambda}$, which defines another family $E_{\lambda}^{m}\rightarrow Y(2)$. In this article we will work with with the product
[…]
We will consider an irreducible curve $\mathcal{C}\subseteq E_{\lambda}^{m}\times E_{\mu}^{n}$, defined over a number field $k$, not contained in a fixed fiber.
Thus, each point $\mathbf{c}\in\mathcal{C}(\mathbb{C})$ defines $m$ points $P_{1}(\mathbf{c}),\ldots,P_{m}(\mathbf{c})$ on the elliptic curve $E_{\lambda(\mathbf{c})}$ and $n$ points $Q_{1}(\mathbf{c}),\ldots,Q_{n}(\mathbf{c})$ on the elliptic curve $E_{\mu(\mathbf{c})}$. We assume that the $P_{i}$’s are linearly independent over $\operatorname{End}(E_{\lambda\rvert_{\mathcal{C}}})$ and the same holds for the $Q_{i}$’s. This is of course equivalent to saying that there are no generic non-trivial linear relations between the $P_{i}$’s and the $Q_{i}$’s. Another way of rephrasing this is to say that $\mathcal{C}$ is not contained in a proper subgroup scheme of $E_{\lambda}^{m}\times E_{\mu}^{n}\rightarrow Y(2)\times Y(2)$.
We define the map
$J:Y(2)$ $\longrightarrow Y(1)=\mathbb{A}^{1}$
$\lambda$ $\longmapsto 2^{8}\dfrac{(\lambda^{2}-\lambda+1)^{3}}{\lambda^{2}(\lambda-1)^{2}}$
which sends $\lambda$ to the $j$-invariant of $E_{\lambda}$. With a slight abuse of notation, we will also denote by $J$ the map $Y(2)^{2}\rightarrow Y(1)^{2}$ obtained by applying $J$ component-wise.
Definition 1.1.
Let $C\subseteq\mathbb{A}^{2}$ be an irreducible curve and let $X,Y$ be the coordinate functions on $\mathbb{A}^{2}$. We say that $C$ is asymmetric (see [Hab10]) if $\deg(X|_{C})\neq\deg(Y|_{C})$.
[…]
We are now ready to state the main result of this article.
Theorem 1.2.
Let $\mathcal{C}\subseteq E_{\lambda}^{m}\times E_{\mu}^{n}$ be an irreducible asymmetric curve defined over $\overline{\mathbb{Q}}$ not contained in a fixed fiber, and define $P_{i},Q_{j}$ as above. Suppose moreover that $E_{\lambda}$ and $E_{\mu}$ are not generically isogenous on $\mathcal{C}$ and that there are no generic non-trivial relations among $P_{1},\ldots,P_{m}$ on $E_{\lambda}$ and among $Q_{1},\ldots,Q_{n}$ on $E_{\mu}$ with coefficients in $\operatorname{End}(E_{\lambda\rvert_{\mathcal{C}}})$ and $\operatorname{End}(E_{\mu\rvert_{\mathcal{C}}})$, respectively. Then, there are at most finitely many $\mathbf{c}\in\mathcal{C}(\mathbb{C})$ such that there exist an isogeny $\phi:E_{\mu(\mathbf{c})}\rightarrow E_{\lambda(\mathbf{c})}$ and $(a_{1},\ldots,a_{m+n})\in\operatorname{End}(E_{\lambda(\mathbf{c})})^{m+n}\setminus\left\{\mathbf{0}\right\}$ with
$a_{1}P_{1}(\mathbf{c})+\ldots+a_{m}P_{m}(\mathbf{c})+a_{m+1}\phi\left(Q_{1}(\mathbf{c})\right)+\ldots+a_{m+n}\phi\left(Q_{n}(\mathbf{c})\right)=O.$
Notice that this theorem is a special case of the Zilber-Pink Conjecture. In particular, combined with results in [BC16], [BC17], [Bar17] and [HP16], it proves the conjecture for an asymmetric curve in $E_{\lambda}^{m}\times E_{\mu}^{n}$ defined over $\overline{\mathbb{Q}}$. For an account on the Zilber-Pink conjecture and other problems of Unlikely Intersections, see [Zan12] and [Pil22].
Notice also that if $E_{\lambda(\mathbf{c}_{0})}$ and $E_{\mu(\mathbf{c}_{0})}$ are isogenous for every $\mathbf{c}_{0}\in\mathcal{C}$, then $\widetilde{\mathcal{C}}$ is not asymmetric (see subsection 2.1), so we could in theory remove this condition from the theorem. However, in view of a possible generalization of this result without the asymmetry condition, we prefer to leave the statement as it is.
Depending on $\pi(\mathcal{C})\subseteq Y(2)^{2}$, we can distinguish three cases:
  1. (i)
    the coordinate functions $\lambda,\mu$ on $\mathcal{C}$ are both non-constant;
  2. (ii)
    (exactly) one between $\lambda$ and $\mu$ is constant and the associated elliptic curve is not CM;
  3. (iii)
    (exactly) one between $\lambda$ and $\mu$ is constant and the associated elliptic curve is CM.
For each $\mathbf{c}\in\mathcal{C}(\mathbb{C})$, let $\rho(\mathbf{c})\in\mathbb{C}$ be such that $\operatorname{End}(E_{\lambda(\mathbf{c})})\cong\mathbb{Z}\left[\rho(\mathbf{c})\right]$.
In case (i), by a theorem by André [And98], there are only finitely many $\mathbf{c}\in\mathcal{C}(\overline{\mathbb{Q}})$ such that $E_{\lambda(\mathbf{c})}$ and $E_{\mu(\mathbf{c})}$ have both complex multiplication. So, recalling that isogenous elliptic curves have the same endomorphism algebra, we can discard those finitely many points and assume that $\rho=0$ and $\mathbf{a}\in\mathbb{Z}^{m+n}\setminus\left\{\mathbf{0}\right\}$.
Similarly, in case (ii), we can assume without loss of generality that $\lambda=\lambda_{0}$ is constant with $E_{\lambda_{0}}$ not CM. Therefore, there are no points $\mathbf{c}\in\mathcal{C}(\overline{\mathbb{Q}})$ such that $E_{\lambda(\mathbf{c})}$ and $E_{\mu(\mathbf{c})}$ have both complex multiplication, so we can take $\rho=0$ and $\mathbf{a}\in\mathbb{Z}^{m+n}\setminus\left\{\mathbf{0}\right\}$ in this case as well.
In case (iii), we can assume again that $\lambda=\lambda_{0}$ is constant. However, in this case there are infinitely many points $\mathbf{c}\in\mathcal{C}(\overline{\mathbb{Q}})$ such that $E_{\lambda(\mathbf{c})}=E_{\lambda_{0}}$ and $E_{\mu(\mathbf{c})}$ are both CM, so we cannot simplify our hypothesis as before. On the other hand, since $\lambda$ is constant, we can choose $\rho$ to be a generator of $\operatorname{End}(E_{\lambda_{0}})\cong\mathbb{Z}[\rho]$.
Our proof of Theorem 1.2 follows the general strategy first introduced by Pila and Zannier in [PZ08] and later used, among the others, by Masser and Zannier [MZ08, MZ10, MZ12] and by Barroero and Capuano [BC16, Bar17, BC17, BC20]. Since the elliptic curves $E_{\lambda}$ and $E_{\mu}$ are analytically isomorphic to the complex tori $\mathbb{C}/\Lambda_{\tau_{1}}$ and $\mathbb{C}/\Lambda_{\tau_{2}}$, where $\Lambda_{\tau}=\mathbb{Z}+\mathbb{Z}\tau$, with $\tau$ in the complex upper half-plane $\mathbb{H}$, we can consider the elliptic logarithms $z_{1}\ldots,z_{m}$ of $P_{1},\ldots,P_{m}$ and $w_{1},\ldots,w_{n}$ of $Q_{1},\ldots,Q_{n}$ and define a uniformization map $(\tau_{1},z_{1}\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n})\mapsto(\lambda,P_{1},\ldots,P_{m},\mu,Q_{1},\ldots,Q_{n})$. By a work of Peterzil and Starchenko, after restricting to a suitable fundamental domain, this map is definable in the o-minimal structure $\mathbb{R}_{\text{an, exp}}$, so the preimage of $\mathcal{C}$ is a definable surface $S$.
Let $\mathcal{C}^{\prime}$ be the subset of $\mathcal{C}$ we want to prove to be finite. Then, the points $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ correspond to points on $S$ lying on subvarieties defined by equations with integer coefficients. We then use a result by Habegger and Pila, which implies that there are $\ll T^{\varepsilon}$ points of $S$ lying on the subvarieties with coefficients bounded in absolute value by $T$, provided that the $z_{i}$ and the $w_{j}$ are algebraically independent over $\mathbb{C}(\tau_{1},\tau_{2})$.
We then use a result by Habegger [Hab10] for asymmetric curves [^1] This is the only step of the proof where we use the assumption on the asymmetry of $\mathcal{C}$, see also remark 5.2., giving height bounds for $\lambda(\mathbf{c}_{0})$, $\mu(\mathbf{c}_{0})$, the $P_{i}(\mathbf{c}_{0})$ and the $Q_{j}(\mathbf{c}_{0})$. By a result of Masser [Mas88], these bounds imply that the coefficients $a_{1},\ldots,a_{m+n}$ of the linear relation between the $m+n$ points
$P_{1}(\mathbf{c}_{0}),\ldots,P_{m}(\mathbf{c}_{0}),\phi(Q_{1}(\mathbf{c}_{0})),\ldots,\phi(Q_{n}(\mathbf{c}_{0}))$
can be taken to be bounded by a positive power of $D_{0}=\left[k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0})):k\right]$. Moreover, all Galois conjugates of $\mathbf{c}_{0}$ are still in $\mathcal{C}^{\prime}$, so that we have at least $D_{0}$ points on $S$ lying on the subvarieties with coefficients bounded in absolute value by some positive power of $D_{0}$. Combining this with the previous bound, we get that $D_{0}$ is bounded and therefore the claim of the theorem, by Northcott’s theorem.
Remark 1.3.
As Gabriel Dill noted, cases (ii) and (iii) can also be deduced from Theorem 1.2 in [Dil21]. Up to considering an isogeny $\Phi:E_{\lambda_{0}}^{m}\times E_{\mu}^{n}\rightarrow E_{\lambda_{0}}^{m}\times E_{\mu}^{n}$ (see also section 2.2 for more details), we can assume that, for some $\ell\geq 0$, $P_{1},\ldots,P_{\ell}$ are constant and $P_{\ell+1},\ldots,P_{m}$ are linearly independent modulo constants over $\operatorname{End}(E_{\lambda_{0}})$. Using the notation from [Dil21], take $A_{0}=E_{\lambda_{0}}^{m-\ell+n}$, $\mathcal{A}=E_{\lambda_{0}}^{m-\ell}\times E_{\mu}^{n}$ and $\Gamma=(\Gamma_{0})^{m-\ell+n}$, where $\Gamma_{0}$ is the divisible hull of the subgroup of $E_{\lambda_{0}}(\overline{\mathbb{Q}})$ generated by $\operatorname{End}(E_{\lambda_{0}})\cdot P_{1},\ldots,\operatorname{End}(E_{\lambda_{0}})\cdot P_{\ell}$. Then, if $\widetilde{\mathcal{C}}$ is the projection of $\Phi(\mathcal{C})$ onto $\mathcal{A}$, $\mathcal{A}_{\Gamma}^{\left[1\right]}\cap\widetilde{\mathcal{C}}$ consists exactly of the points described in Theorem 1.2 and, by [Dil21, Theorem 1.2], we get that either this intersection is finite or that the generic point $\mathcal{C}_{\xi}\in\widetilde{\mathcal{C}}$ is contained in the translate of a proper abelian subvariety of $\mathcal{A}_{\xi}$ by a point in
$(\mathcal{A}_{\xi})_{\mathrm{tors}}+\mathrm{Tr}(\mathcal{A}_{\xi})=E_{\lambda_{0}}^{m-\ell}\times(E_{\mu}^{n})_{\mathrm{tors}}.$
However, the latter means that either $Q_{1},\ldots,Q_{n}$ are generically linearly dependent or that there is a non trivial linear relation modulo constants involving $P_{1},\ldots,P_{\ell}$ and $P_{\ell+1},\ldots,P_{m}$, contradicting our assumptions.
2. Preliminaries
2.1. Isogenies and modular curves
Let $E_{1}\cong\mathbb{C}/\Lambda_{1}$ and $E_{2}\cong\mathbb{C}/\Lambda_{2}$ be two elliptic curves defined over $\mathbb{C}$. Up to homothety, we can choose $\Lambda_{1}=\mathbb{Z}+\mathbb{Z}\tau_{1}$ and $\Lambda_{1}=\mathbb{Z}+\mathbb{Z}\tau_{2}$, for some $\tau_{1},\tau_{2}$ in the upper half-plane $\mathbb{H}$.
Recall that for each isogeny $\phi:E_{1}\rightarrow E_{2}$ there exists a unique non-zero complex number $\alpha$ such that $\alpha\Lambda_{1}\subseteq\Lambda_{2}$ and $\phi$ corresponds to the multiplication by $\alpha$ map $\mathbb{C}/\Lambda_{1}\rightarrow\mathbb{C}/\Lambda_{2}$.
Therefore, if $E_{1}$ and $E_{2}$ are isogenous, then there exists $\alpha\in\mathbb{C}\setminus\left\{0\right\}$ and integers $A,B,C,D$ not all zero (not necessarily coprime) such that
$\alpha\cdot\tau_{1}$ $=A\tau_{2}+B$
$\alpha\cdot 1$ $=C\tau_{2}+D$
thus
$\tau_{1}=\dfrac{A\tau_{2}+B}{C\tau_{2}+D}.$
Moreover, the converse is also true. If $\tau_{1},\tau_{2}\in\mathbb{H}$ and $\tau_{1}=\frac{A\tau_{2}+B}{C\tau_{2}+D}$ for integers $A,B,C,D$, then there exists an isogeny $\phi:E_{1}\rightarrow E_{2}$ corresponding to $\alpha=C\tau_{2}+D$.
More generally, we have an action of the group $\operatorname{GL}_{2}^{+}(\mathbb{Q})$ (here $+$ means that the matrices have positive determinant) on the upper half-plane $\mathbb{H}$ which is given by
$M\tau=\dfrac{a\tau+b}{c\tau+d}$
for $M=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{GL}_{2}^{+}(\mathbb{Q})$. If $M\in\textrm{Mat}(\mathbb{Z},2)$, we say that $M$ is primitive if $\gcd(a,b,c,d)=1$.
We say that an isogeny $\phi$ is cyclic if $\ker\phi$ is a (finite) cyclic group. Then it is known (see [DS05, Section 1.3]) that any isogeny can be written as the composition of a cyclic isogeny and a multiplication-by-$n$ isogeny, for some integer $n$. In particular, cyclic isogenies $E_{1}=\mathbb{C}/\Lambda_{1}\rightarrow E_{2}=\mathbb{C}/\Lambda_{2}$ correspond to relations $\tau_{1}=M\tau_{2}$ with $M$ primitive. In this case, the degree of the isogeny is equal to $\det M$.
The following theorem provides an effective bound for the size of the integers $A,B,C,D$ when the degree of the isogeny is fixed.
Lemma 2.1 ([HP12], Lemma 5.2.).
There exists an absolute constant $c>0$ with the following property: if $E_{1},E_{2}$ are elliptic curves defined over $\mathbb{C}$ and there exists a cyclic isogeny $\phi:E_{1}\rightarrow E_{2}$ of degree $N$, then there are integers $A,B,C,D$ such that
$AD-BC=N\qquad\tau_{1}=\dfrac{A\tau_{2}+B}{C\tau_{2}+D}\qquad\left|A\right|,\left|B\right|,\left|C\right|,\left|D\right|\leq cN^{10}.$
Finally, recall that the modular polynomials $\Phi_{N}(X,Y)\in\mathbb{Z}\left[X,Y\right]$ are the irreducible symmetric polynomials parametrizing pairs of isomorphism classes of elliptic curves with a cyclic isogeny of degree $N$ between them [Lan87, Chapter 5]. In other words, $\Phi_{N}(j_{1},j_{2})=0$ if and only if there exists a cyclic isogeny of degree $N$ between the elliptic curves with $j$-invariants $j_{1}$ and $j_{2}$. We then define the classical modular curve $Y_{0}(N)\subset Y(1)^{2}=\mathbb{A}^{2}$ as the plane curve defined by the equation $\Phi_{N}(X,Y)=0$.
2.2. Uniformization
Let $\mathcal{A}$ be the quasi-projective variety in $Y(2)\times(\mathbb{P}^{2})^{m}\times Y(2)\times(\mathbb{P}^{2})^{n}$ with coordinates
$(\lambda,\left[X_{1}:Y_{1}:Z_{1}\right],\ldots,\left[X_{m}:Y_{m}:Z_{m}\right],\mu,\left[U_{1}:V_{1}:W_{1}\right],\ldots,\left[U_{n}:V_{n}:W_{n}\right])$
and defined by the $n+m$ equations
$Y_{i}^{2}Z_{i}$ $=X_{i}(X_{i}-Z_{i})(X_{i}-\lambda Z_{i})$ $i=1,\ldots,m$
$V_{j}^{2}W_{j}$ $=U_{j}(U_{j}-W_{j})(U_{j}-\mu W_{j})$ $j=1,\ldots,n.$
We set $P_{i}=\left[X_{i}:Y_{i}:Z_{i}\right]$ and $Q_{j}=\left[U_{j}:V_{j}:W_{j}\right]$ and we have an irreducible curve $\mathcal{C}\subseteq\mathcal{A}$ defined over a number field $k$ such that the projection of $\mathcal{A}$ to $Y(2)\times Y(2)$ restricts to rational functions $\lambda$ and $\mu$ on $\mathcal{C}$ not both constant.
The aim of this section is to define a uniformization map for $\mathcal{A}$.
As said before, any elliptic curve over $\mathbb{C}$ is analytically isomorphic to a complex torus $\mathbb{C}/\Lambda_{\tau}$, where $\tau$ has positive imaginary part and $\Lambda_{\tau}$ is the lattice generated by $1$ and $\tau$, with fundamental domain
$\mathcal{L}_{\tau}=\left\{z\in\mathbb{C}:z=x+\tau y,x,y\in\left[0,1\right)\right\}.$
The classical Weierstrass Weierstrass-$\wp$-function $\wp(z,\Lambda_{\tau})=\wp(z,\tau)$ associated to the lattice $\Lambda_{\tau}$, is $\Lambda_{\tau}$-periodic and satisfies the following differential equation
$(\wp(z,\tau)^{\prime})^{2}=4\wp(z,\tau)^{3}-g_{2}(\tau)\wp(z,\tau)-g_{3}(\tau)$
where $\wp(z,\tau)^{\prime}=\frac{d}{dz}\wp(z,\tau)$. Then, the zeros of the polynomial $4X^{3}-g_{2}(\tau)X-g_{3}(\tau)$ are exactly the values of Weierstrass-$\wp$ at the half-periods:
$e_{1}(\tau)=\wp\left(\dfrac{1}{2},\tau\right)\quad e_{2}(\tau)=\wp\left(\dfrac{1+\tau}{2},\tau\right)\quad e_{3}(\tau)=\wp\left(\dfrac{\tau}{2},\tau\right).$
Note that the $e_{i}(\tau)$ are pairwise distinct and that $e_{3}(\tau)-e_{1}(\tau)$ has a regular square root for all $\tau\in\mathbb{H}$. Therefore, we can define
$\xi(z,\tau)=\dfrac{\wp(z,\tau)-e_{1}(\tau)}{e_{3}(\tau)-e_{1}(\tau)}\quad\text{and}\quad\eta(z,\tau)=\dfrac{\wp(z,\tau)^{\prime}}{2(e_{3}(\tau)-e_{1}(\tau))^{\frac{3}{2}}}$
so that we have the following relation
$\eta(z,\tau)^{2}=\xi(z,\tau)(\xi(z,\tau)-1)(\xi(z,\tau)-L(\tau))$
where
$L(\tau)=\dfrac{e_{2}(\tau)-e_{1}(\tau)}{e_{3}(\tau)-e_{1}(\tau)}.$ (2.1)
This gives a parametrization of the Legendre family via the map $(z,\tau)\mapsto\left(L(\tau),P(z,\tau)\right)$, where
$P(z,\tau)=\begin{cases}\left[\xi(z,\tau):\eta(z,\tau):1\right]&\text{ if }z\not\in\Lambda_{\tau}\\ \left[0:1:0\right]&\text{ otherwise }\end{cases}$
Finally, define the map $\varphi:\mathbb{H}\times\mathbb{C}^{m}\times\mathbb{H}\times\mathbb{C}^{n}\rightarrow\mathcal{A}(\mathbb{C})$ sending $\left(\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}\right)$ to $\left(L(\tau_{1}),P(z_{1},\tau_{1}),\ldots,P(z_{m},\tau_{1}),L(\tau_{2}),P(w_{1},\tau_{2}),\ldots,P(w_{n},\tau_{2})\right)$. Since this map is not injective we would like to find a subset of the domain over which it is possible to define a univalued inverse function of $\varphi$.
By [For51, Sec. 70], there exists a finite index subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{Z})$ such that $L(\gamma\tau)=L(\tau)$ for every $\gamma\in\Gamma$. Moreover, as a fundamental domain for the action of $\Gamma$ on $\mathbb{H}$ one can take the union of six suitably chosen fundamental domains for the action of $\mathrm{SL}_{2}(\mathbb{Z})$ (see [For51, Fig. 48 and 49]). We will call this set $\mathcal{B}$ and define
=$\mathcal{F}_{\mathcal{B}}=\left\{\left(\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}\right):\tau_{1},\tau_{2}\in\mathcal{B},z_{1},\ldots,z_{m}\in\mathcal{L}_{\tau_{1}},w_{1},\ldots,w_{n}\in\mathcal{L}_{\tau_{2}}\right\}.$
Then, $\varphi$ has a univalued inverse $\mathcal{A}(\mathbb{C})\rightarrow\mathcal{F}_{\mathcal{B}}$ and we set
$\mathcal{Z}=\varphi^{-1}(\mathcal{C}(\mathbb{C}))\cap\mathcal{F}_{\mathcal{B}}.$ (2.2)
In general, let $C\subseteq E_{\lambda}^{m}\times E_{\mu}^{n}$ be an irreducible curve not contained in a fixed fiber, and define $P_{i},Q_{j}$ and $\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}$ as before. Assume also that $E_{\lambda}$ and $E_{\mu}$ are not generically isogenous on $C$ and that there are no generic non-trivial relations among $P_{1},\ldots,P_{m}$ on $E_{\lambda}$ and among $Q_{1},\ldots,Q_{n}$ on $E_{\mu}$ with coefficients in $\operatorname{End}(E_{\lambda})$ and $\operatorname{End}(E_{\mu})$, respectively.
Then, take a small open disc $D$ on $C$ and consider $\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}$ as holomorphic functions on $D$.
Furthermore, in case (i) let $\ell=0$, while in case (ii) and (iii) let $\ell\geq 0$ be the greatest integer such that there are $\widetilde{a}_{i,j}\in\operatorname{End}(E_{\lambda_{0}})$ and $\widetilde{P}_{j}\in E_{\lambda_{0}}(\mathbb{C})$, $i=1,\ldots,m$ and $j=1,\ldots,\ell$, such that the vectors $(\widetilde{a}_{1,j},\ldots,\widetilde{a}_{m,j})$, for $j=1,\ldots,\ell$, are $\mathbb{Z}$-linearly independent and
$\widetilde{a}_{1,j}P_{1}+\ldots+\widetilde{a}_{m,j}P_{m}=\widetilde{P}_{j}.$
Hence, we can consider the isogeny
$\Phi:E_{\lambda_{0}}^{m}\times E_{\mu}^{n}$ $\longrightarrow E_{\lambda_{0}}^{m}\times E_{\mu}^{n}$
$(P_{1},\ldots,P_{m},Q_{1},\ldots,Q_{n})$ $\longmapsto\left(\sum\limits_{i=1}^{m}\widetilde{a}_{i,1}P_{i},\ldots,\sum\limits_{i=1}^{m}\widetilde{a}_{i,\ell}P_{i},P_{\ell+1},\ldots,P_{m},Q_{1},\ldots,Q_{n}\right)$
which sends $(\mathbf{P},\mathbf{Q})\in C$ to $(\widetilde{P}_{1},\ldots,\widetilde{P_{\ell}},P_{\ell+1},\ldots,P_{m},\mathbf{Q})\in\Phi(C)$.
Under these assumptions, we have the following transcendence result.
Lemma 2.2.
$z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}$ are algebraically independent over $\mathbb{C}(\tau_{1},\tau_{2})$.
Proof.
In case (i) we can apply Corollary 2.5 from [BC17] (which is based on a result by Bertrand [Ber09]).
In case (ii) and (iii), let $C^{\prime}=\Phi(C)$ and notice that $C^{\prime}\subseteq\left\{\left(\widetilde{P}_{1},\ldots,\widetilde{P_{\ell}}\right)\right\}\times E_{\lambda_{0}}^{m-\ell}\times E_{\mu}^{n}$.
Let also $F=\mathbb{C}\left(\tau_{2}\right)$ and assume by contradiction that
$\mathrm{tr.deg}_{F}F\left(z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}\right)<m+n-\ell.$
Then, if $\widetilde{z}_{1},\ldots,\widetilde{z}_{\ell}$ are elliptic logarithms of the (constant) points $\widetilde{P}_{1},\ldots,\widetilde{P}_{\ell}$, on $C^{\prime}$ we must have
$\mathrm{tr.deg}_{F}F\left(\widetilde{z}_{1},\ldots,\widetilde{z}_{\ell},z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}\right)=\mathrm{tr.deg}_{F}F\left(z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}\right)<m+n-\ell.$
Applying Theorem 7.1 from [Dil21] to $C^{\prime}$, we deduce the existence of a subvariety $\mathcal{W}\subseteq E_{\lambda_{0}}^{m}\times E_{\mu}^{n}$ which is a translate of an abelian subscheme of $E_{\lambda_{0}}^{m}\times E_{\mu}^{n}$ by a point in $E_{\lambda_{0}}^{m}(\overline{\mathbb{Q}})\times(E_{\mu}^{n})_{\mathrm{tors}}$, containing $C^{\prime}$ and such that $\dim(\mathcal{W})\leq m+n-\ell$.
Since $Q_{1},\ldots,Q_{n}$ are linearly independent by hypothesis, this implies that there are $\widetilde{a}_{i,\ell+1}\in\operatorname{End}(E_{\lambda_{0}})$ and $\widetilde{P}_{\ell+1}\in E_{\lambda_{0}}(\overline{\mathbb{Q}})$, $i=\ell+1,\ldots,m$, such that $\widetilde{a}_{\ell+1,\ell+1},\ldots,\widetilde{a}_{m,\ell+1}$ are not all zero and
$\widetilde{a}_{\ell+1,\ell+1}P_{\ell+1}+\ldots+\widetilde{a}_{m,\ell+1}P_{m}=\widetilde{P}_{\ell+1}$
contradicting the maximality of $\ell$ and proving that
$\mathrm{tr.deg}_{F}F\left(z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}\right)=m+n-\ell.$
2.3. Heights
Let $h$ denote the logarithmic absolute Weil height on $\mathbb{P}^{N}$, as defined in [BG06] and, if $\alpha$ is an algebraic number, define $h(\alpha)=h\left(\left[1:\alpha\right]\right)$. Define also the multiplicative Weil height as $H(P)=\exp(h(P))$.
For an elliptic curve $E$ defined over $\overline{\mathbb{Q}}$ and a point $P\in E(\overline{\mathbb{Q}})\subseteq\mathbb{P}^{2}(\overline{\mathbb{Q}})$, we also have the Néron-Tate height $\widehat{h}$, defined as follows (see also [Sil09, VII.9]):
$\widehat{h}(P)=\lim\limits_{n\rightarrow\infty}\dfrac{1}{4^{n}}h(2^{n}P)$
Proposition 2.3.
Let $E_{1},E_{2}$ two elliptic curves defined over $\overline{\mathbb{Q}}$ and let $\phi:E_{1}\rightarrow E_{2}$ be an isogeny also defined over $\overline{\mathbb{Q}}$. Denote by $\widehat{h}_{1}$ and $\widehat{h}_{2}$ the Néron-Tate heights on $E_{1}$ and $E_{2}$, respectively. Then for any $P\in E_{1}(\overline{\mathbb{Q}})$, we have
$\widehat{h}_{2}(\phi(P))=\deg\phi\cdot\widehat{h}_{1}(P).$
Proof.
Recall that the height $h$, given by the embedding of an elliptic curve $E$ into $\mathbb{P}^{2}$, is exactly the height associated to the divisor $3(O)$ as described in [HS13, Section B.3], where $O$ is the identity element on $E$. Moreover, if $O_{i}$ is the identity on $E_{i}$ and $\phi:E_{1}\rightarrow E_{2}$ is an isogeny, we have
$\phi^{*}2(O_{2})=2\sum\limits_{T\in\ker\phi}(T)\sim 2(T^{\prime})+(2\deg\phi-2)(O_{1})$
where $T^{\prime}=\sum\limits_{T\in\ker\phi}T$ is either $O_{1}$ or a non trivial 2-torsion point. In either case, $2(T^{\prime})\sim 2(O_{1})$, so that the pull-back $\phi^{*}2(O_{2})$ is linearly equivalent to $(2\deg\phi)(O_{1})$. Then, using [HS13, Theorem B.5.6], we get
$2\widehat{h}_{2}(\phi(P))$ $=2\widehat{h}_{E_{2},3(O_{2})}(\phi(P))=\widehat{h}_{E_{2},6(O_{2})}(\phi(P))$
$=3\widehat{h}_{E_{2},2(O_{2})}(\phi(P))=3\widehat{h}_{E_{1},\phi^{*}2(O_{2})}(P)=$
$=3\widehat{h}_{E_{1},(2\deg\phi)(O_{1})}(P)=2\deg\phi\cdot\widehat{h}_{E_{1},3(O_{1})}(P)=2\deg\phi\cdot\widehat{h}_{1}(P)$
which is equivalent to $\widehat{h}_{2}(\phi(P))=\deg\phi\cdot\widehat{h}_{1}(P)$. ∎
Using the same notation as in the previous section, we have that if $\mathbf{c}\in\mathcal{C}\left(\overline{\mathbb{Q}}\right)$, then the standard properties of heights imply that, if $\lambda$ and $\mu$ are both non-constant, we have
$h(P_{i}(\mathbf{c}))\ll h(\lambda(\mathbf{c}))+1\quad\textrm{and}\quad h(Q_{j}(\mathbf{c}))\ll h(\mu(\mathbf{c}))+1$
for every $i=1,\ldots,m$ and $j=1,\ldots,n$. In case (ii) and (iii), if $\lambda=\lambda_{0}$ is constant, we have that $h(P_{i}(\mathbf{c}))\ll h(\mu(\mathbf{c}))+1$, as we can use $\mu$ as uniformizing parameter on the base $\pi(\mathcal{C})=\left\{\lambda_{0}\right\}\times\mathbb{A}^{1}$. Moreover, note that if $\mathcal{C}$ is defined over a number field $k$, we also have
$\left[k(\mathbf{c}):k\right]\ll\left[k(\lambda(\mathbf{c}),\mu(\mathbf{c})):k\right].$
Finally, we will also need another definition of height (from [Pil09]).
Definition 2.4.
If $\alpha$ is a complex number, we define
$H_{1}(\alpha):=\begin{cases}H(\alpha)=\max\left\{\left|p\right|,\left|q\right|\right\}\quad\text{ if }\alpha=\frac{p}{q}\in\mathbb{Q}\\ +\infty\quad\text{otherwise}\end{cases}$
For $(\alpha_{1},\ldots,\alpha_{N})\in\mathbb{C}^{N}$, we also define $H_{1}(\alpha_{1},\ldots,\alpha_{N})=\max\left\{H_{1}(\alpha_{i})\right\}$.
2.4. Complex Multiplication
Given a $\lambda_{0}\in Y(2)$ such that $E_{\lambda_{0}}$ has complex multiplication, we know that the associated $\tau_{0}\in\mathcal{B}$ is an algebraic number of degree 2, with minimal polynomial $aX^{2}+bX+c$ and discriminant $\Delta_{0}=b^{2}-4ac<0$. In this case, we know by [Lan87, Theorem 1, p. 90], that
$\operatorname{End}(E_{\lambda_{0}})=\mathcal{O}_{\lambda_{0}}=\mathbb{Z}\left[\rho_{0}\right]$
where $\rho_{0}=\frac{\Delta_{0}+\sqrt{\Delta_{0}}}{2}$.
By Theorem II.4.3. of [Sil94],
$\left[\mathbb{Q}(j_{0}):\mathbb{Q}\right]=cl(\mathcal{O}_{\lambda_{0}})$
where $j_{0}$ is the $j$-invariant of $E_{\lambda_{0}}$ (which is algebraic by [Sil94, Proposition II.2.1]) and $cl(\mathcal{O}_{\lambda_{0}})$ is the class number of $\mathcal{O}_{\lambda_{0}}$.
Moreover, a theorem of Siegel in the form of Theorem 1.2 of [Bre01] gives us the estimate
$\left|\Delta_{0}\right|^{\frac{1}{2}-\epsilon}\ll_{\epsilon}cl(\mathcal{O}_{\lambda_{0}})\ll_{\epsilon}\left|\Delta_{0}\right|^{\frac{1}{2}+\epsilon}$
so that, in particular, we have $\left|\Delta_{0}\right|\ll\left[\mathbb{Q}(j_{0}):\mathbb{Q}\right]^{3}$. Finally, using Proposition 2.3 and the fact that the endomorphism $\rho_{0}$ has degree $\left(\Delta_{0}^{2}-\Delta_{0}\right)/4$, we get that
$\widehat{h}(\rho_{0}P)\ll\left|\Delta_{0}\right|^{2}\widehat{h}(P)\ll\left[\mathbb{Q}(j_{0}):\mathbb{Q}\right]^{6}\widehat{h}(P)\ll\left[\mathbb{Q}(\lambda_{0}):\mathbb{Q}\right]^{6}\widehat{h}(P)$ (2.3)
for every $P\in E_{\lambda_{0}}\!\!\left(\overline{\mathbb{Q}}\right)$.
3. O-minimality and definable sets
In this section we recall the basic properties and some results about o-minimal structures. For more details see [vdD98] and [vdDM96].
Definition 3.1.
A structure is a sequence $\mathcal{S}=\left(\mathcal{S}_{N}\right)$, $N\geq 1$, where each $\mathcal{S}_{N}$ is a collection of subsets of $\mathbb{R}^{N}$ such that, for each $N,M\geq 1$:
  • $\mathcal{S}_{N}$ is a boolean algebra (under the usual set-theoretic operations);
  • $\mathcal{S}_{N}$ contains every semi-algebraic subset of $\mathbb{R}^{N}$;
  • if $A\in\mathcal{S}_{N}$ and $B\in\mathcal{S}_{M}$, then $A\times B\in\mathcal{S}_{N+M}$;
  • if $A\in\mathcal{S}_{N+M}$, then $\pi(A)\in\mathcal{S}_{N}$, where $\pi:\mathbb{R}^{N+M}\rightarrow\mathbb{R}^{N}$ is the projection onto the first $N$ coordinates.
If $\mathcal{S}$ is a structure and, in addition,
  • $\mathcal{S}_{1}$ consists of all finite union of open intervals and points
then $\mathcal{S}$ is called an o-minimal structure.
Given a structure $\mathcal{S}$, we say that $S\subseteq\mathbb{R}^{N}$ is a definable set if $S\in\mathcal{S}_{N}$.
Given $S\subseteq\mathbb{R}^{N}$ and a function $f:S\rightarrow\mathbb{R}^{M}$, we say that $f$ is a definable function if its graph $\left\{(x,y)\in\mathbb{R}^{N}\times\mathbb{R}^{M}:x\in S,y=f(x)\right\}$ is a definable set. One can easily prove that images and preimages of definable sets via definable functions are still definable.
Let $U\subseteq\mathbb{R}^{N+M}$. For $t_{0}\in\mathbb{R}^{M}$, we set $U_{t_{0}}=\left\{x\in\mathbb{R}^{N}:(t_{0},x)\in U\right\}$ and call $U$ a family of subsets of $\mathbb{R}^{N}$, while $U_{t_{0}}$ is called the fiber of $U$ above $t_{0}$. If $U$ is a definable set, then we call it a definable family and it is easy to prove that the fibers $U_{t_{0}}$ are also definable.
Proposition 3.2 ([vdDM96], 4.4).
Let $U$ be a definable family in a fixed o-minimal structure $\mathcal{S}$. Then, there exists an integer $n$ such that each fiber of $U$ has at most $n$ connected components.
While there are many examples of o-minimal structures (see [vdDM96]), in this article we will work with the structure $\mathbb{R}_{\text{an,exp}}$, which was proved to be o-minimal by van den Dries and Miller [vdDM94].
For a family $Z\subseteq\mathbb{R}^{M}\times\mathbb{R}^{N}=\mathbb{R}^{M+N}$ and a positive real number $T$ define
$Z^{\sim}(\mathbb{Q},T):=\left\{(y,z)\in Z:y\in\mathbb{Q}^{M},H_{1}(y)\leq T\right\}$
where $H_{1}(y)$ is the 1-polynomial height defined in the previous section and let $\pi_{1},\pi_{2}$ be the projections of $Z$ to the first $M$ and last $N$ coordinates, respectively.
Proposition 3.3 ([HP16], Corollary 7.2).
Let $Z\subseteq\mathbb{R}^{M+N}$ be a definable set. For every $\varepsilon>0$ there exists a positive constant $c=c(Z,\varepsilon)$ with the following property. If $T\geq 1$ and $\lvert\pi_{2}(Z^{\sim}(\mathbb{Q},T))\rvert>cT^{\varepsilon}$, then there exists a continuous definable function $\delta:\left[0,1\right]\rightarrow Z$ such that:
  1. (1)
    the composition $\pi_{1}\circ\delta:\left[0,1\right]\rightarrow\mathbb{R}^{M}$ is semi-algebraic and its restriction to $(0,1)$ is real analytic;
  2. (2)
    the composition $\pi_{2}\circ\delta:\left[0,1\right]\rightarrow\mathbb{R}^{N}$ is non-constant.
Lastly, we want to prove that the set $\mathcal{Z}$ defined in (2.2) is definable in $\mathbb{R}_{\mathrm{an,exp}}$. In the following, definability will always be considered in $\mathbb{R}_{\mathrm{an,exp}}$, and we say that $X\subseteq\mathbb{C}^{N}$ is definable if the set $\left\{\left(\mathrm{Re}(z_{1}),\mathrm{Im}(z_{1}),\ldots,\mathrm{Re}(z_{N}),\mathrm{Im}(z_{N})\right):(z_{1},\ldots,z_{N})\in X\right\}\subseteq\mathbb{R}^{2N}$ is definable. Similarly, a function $f:X\rightarrow\mathbb{C}$ is definable if and only if $\mathrm{Re}(f)$ and $\mathrm{Im}(f)$ are both definable.
Let $\mathcal{D}$ be the usual fundamental domain for the action of $\mathrm{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$, then the restriction of $\wp(z,\tau)$ to $\left\{(z,\tau):\tau\in\mathcal{D},z\in\mathcal{L}_{\tau}\right\}$ is definable by work of Peterzil and Starchenko [PS05]. Therefore, $\wp(z,\tau)$ is definable even if restricted to $\left\{(z,\tau):\tau\in\gamma\mathcal{D},z\in\mathcal{L}_{\tau}\right\}$, for any fundamental domain $\gamma\mathcal{D}$ for $\mathrm{SL}_{2}(\mathbb{Z})$. Since $\mathcal{B}$ is the union of six such fundamental domains, we have that $\wp(z,\tau)$ is also definable when restricted to $\left\{(z,\tau):\tau\in\mathcal{B},z\in\mathcal{L}_{\tau}\right\}$. Thus, the uniformization map $\varphi$, defined in the previous section and restricted to $\mathcal{F}_{\mathcal{B}}$, is definable. Since $\mathcal{C}$ is semi-algebraic and $\mathcal{F}_{\mathcal{B}}$ is definable, we get that $\mathcal{Z}=\varphi^{-1}(\mathcal{C})\cap\mathcal{F}_{\mathcal{B}}$ is definable.
4. The main estimate
For every $T\geq 1$ define the set
[…]
$\tau_{2}=\dfrac{A\tau_{1}+B}{C\tau_{1}+D},\exists\left(a_{1},\ldots,a_{m+n},b_{1},\ldots,b_{m+n}\right)\in\mathbb{Z}^{2m+2n}\setminus\left\{\mathbf{0}\right\}\text{ with }\max\lvert a_{i}\rvert,\left|b_{i}\right|\leq T$
$\text{and }\sum\limits_{i=1}^{m}(a_{i}+b_{i}\rho)z_{i}+(C\tau_{1}+D)\cdot\sum\limits_{j=1}^{n}(a_{m+j}+b_{m+j}\rho)w_{j}\in\mathbb{Z}+\mathbb{Z}\tau_{1}\Bigg{\}}$
where $\mathcal{Z}$ is the set defined in (2.2) and $\rho$ is either 00 in case (i) and (ii), or a fixed quadratic integer in case (iii).
The goal of this section is to prove the following result.
Proposition 4.1.
Under the hypotheses of Theorem 1.2, for all $\varepsilon>0$, we have #$\#\mathcal{Z}(T)\ll_{\varepsilon}T^{\varepsilon}$, for all $T\geq 1$.
To prove this, we will apply Proposition 3.3 to the definable set $W$ consisting of tuples of the form
$(\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n},A,B,C,D,\gamma_{1},\gamma_{2},\zeta_{1},\theta_{1},x_{1},y_{1},\ldots,x_{m},y_{m},\zeta_{2},\theta_{2},u_{1},v_{1},\ldots,u_{n},v_{n})$
in $\mathbb{R}^{2m+2n+6}\times\mathbb{R}^{2m+2n+4}$, satisfying the following relations:
$\left(\alpha_{1},\ldots,\beta_{m+n}\right)\neq\mathbf{0}\qquad\qquad\left(C(\zeta_{1}+\theta_{1}I)+D\right)(\zeta_{2}+\theta_{2}I)=A(\zeta_{1}+\theta_{1}I)+B$
$\left(\zeta_{1}+\theta_{1}I,x_{1}+y_{1}I,\ldots,x_{m}+y_{m}I,\zeta_{2}+\theta_{2}I,u_{1}+v_{1}I,\ldots,u_{n}+v_{n}I\right)\in\mathcal{Z}$
$\sum\limits_{i=1}^{m}(\alpha_{i}+\beta_{i}\rho)(x_{i}+y_{i}I)+\left(C(\zeta_{1}+\theta_{1}I)+D\right)\sum\limits_{j=1}^{n}(\alpha_{m+j}+\beta_{m+j}\rho)(u_{j}+v_{j}I)=\gamma_{1}+\gamma_{2}(\zeta_{1}+\theta_{1}I)$
where $I$ is the imaginary unit. In particular, we consider for each $T\geq 1$
$W^{\sim}(\mathbb{Q},T):=\left\{(\alpha_{1},\ldots,v_{n})\in W:H_{1}(\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n},A,B,C,D,\gamma_{1},\gamma_{2})\leq T\right\}$
where we recall that $H_{1}(\alpha_{1},\ldots,\gamma_{2})$ is finite if and only if $(\alpha_{1},\ldots,\gamma_{2})\in\mathbb{Q}^{2m+2n+6}$.
Let $\pi_{1},\pi_{2}$ be the projections on the first $62m+2n+$ 6 and the last $42m+2n+$ 4 coordinates, respectively.
Lemma 4.2.
For every $\varepsilon>0$, #$\#\pi_{2}\left(W^{\sim}(\mathbb{Q},T)\right)\ll_{\varepsilon}T^{\varepsilon}$, for all $T\geq 1$.
Proof.
Fix $\varepsilon>0$ and let $c=c(W,\varepsilon)$ be the constant given by Proposition 3.3. Suppose also that #$\#\pi_{2}\left(W^{\sim}(\mathbb{Q},T)\right)>cT^{\varepsilon}$ for some $T\geq 1$.
Then, by Proposition 3.3, there exists a continuous definable function $\delta:\left[0,1\right]\rightarrow W$ such that $\delta_{1}=\pi_{1}\circ\delta:\left[0,1\right]\rightarrow\mathbb{R}^{2m+2n+6}$ is semi-algebraic and $\delta_{2}=\pi_{2}\circ\delta:\left[0,1\right]\rightarrow\mathbb{R}^{2m+2n+4}$ is non-constant. Thus, there exists an infinite connected $J\subseteq\left[0,1\right]$ such that $\delta_{1}(J)$ is an algebraic curve arc and $\delta_{2}(J)$ has positive dimension.
Consider the coordinates
$\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n},A,B,C,D,\gamma_{1},\gamma_{2},\zeta_{1},\theta_{1},x_{1},y_{1},\ldots,x_{m},y_{m},\zeta_{2},\theta_{2},u_{1},v_{1},\ldots,u_{n},v_{n}$
as functions on $J$ and define
$\tau_{1,2}=\zeta_{1,2}+\theta_{1,2}I,\qquad z_{p}=x_{p}+y_{p}I,\qquad w_{q}=u_{q}+v_{q}I$
with $p=1,\ldots,m$ and $q=1,\ldots n$.
On $J$, the functions $\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n},A,B,C,D,\gamma_{1},\gamma_{2},$ satisfy $52m+2n+6-1=2m+2n+$ 5 independent algebraic relations over $\mathbb{C}$ (because they are functions on a curve). Since $\left(\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n}\right)\neq\mathbf{0}$ and by the relations
$\sum\limits_{p=1}^{m}(\alpha_{p}+\beta_{p}\rho)z_{p}+(C\tau_{1}+D)\cdot\sum\limits_{q=1}^{n}(\alpha_{m+q}+\beta_{m+q}\rho)w_{q}=\gamma_{1}+\gamma_{2}\tau_{1}$
$\left(C\tau_{1}+D\right)\tau_{2}=A\tau_{1}+B$
$\widetilde{a}_{1,j}P_{1}+\ldots+\widetilde{a}_{m,j}P_{m}=\widetilde{P}_{j}$
for $j=1,\ldots,\ell$, it follows that the $62m+2n+6+m+n=3m+3n+$ 6 functions
$\alpha_{1},\ldots,\alpha_{m+n},\beta_{1},\ldots,\beta_{m+n},A,B,C,D,\gamma_{1},\gamma_{2},z_{1},\ldots,z_{m},w_{1},\ldots,w_{n}$
satisfy $72m+2n+5+2+\ell=2m+2n+\ell+$ 7 independent algebraic relations over $F=\mathbb{C}\left(\tau_{1},\tau_{2}\right)$. Therefore,
$\mathrm{trdeg}_{F}F\left(z_{1},\ldots,z_{m},w_{1},\ldots,w_{n}\right)\leq 3m+3n+6-(2m+2n+\ell+7)=m+n-\ell-1$
which implies that $z_{\ell+1},\ldots,z_{m},w_{1},\ldots,w_{n}$ are algebraically dependent over $F$ and thus, by Lemma 2.2, we get a contradiction, proving the proposition. ∎
Lemma 4.3.
There exists a positive constant $c^{\prime}=c^{\prime}(\mathcal{Z})$ such that for all $(z_{1},,\ldots,z_{m},w_{1},\ldots,w_{n})\in\mathbb{C}^{m+n}$ and for all $T\geq 1$, there are at most $c^{\prime}$ pairs $(\tau_{1},\tau_{2})\in\mathbb{C}^{2}$ with
$\left(\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}\right)\in\mathcal{Z}(T)$
Proof.
Let
$\widetilde{\pi}:\mathcal{Z}$ $\longrightarrow\mathbb{C}^{m+n}$
$\left(\tau_{1},\mathbf{z},\tau_{2},\mathbf{w}\right)$ $\longmapsto(\mathbf{z},\mathbf{w})$
By o-minimality, if $\widetilde{\pi}^{-1}(\mathbf{z},\mathbf{w})$ has dimension 0, then by Proposition 3.2 there is a uniform bound on its cardinality, which depends only on $\mathcal{Z}$. Hence, we only need to prove that if $(\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n})\in\mathcal{Z}(T)$ for some $T$, then $\widetilde{\pi}^{-1}(\mathbf{z},\mathbf{w})$ has dimension 0.
Suppose that it has positive dimension, then $\left(z_{1},\ldots,z_{m}\right)$ and $\left(w_{1},\ldots,w_{n}\right)$ are algebraically dependent over $\mathbb{C}(\tau_{1},\tau_{2})$ and therefore $\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n}$ are algebraically dependent on an open disc in $\mathcal{C}$, which contradicts Lemma 2.2. ∎
Proof of Proposition 4.1.
If $(\tau_{1},z_{1},\ldots,z_{m},\tau_{2},w_{1},\ldots,w_{n})\in\mathcal{Z}(T)$, then there are integers $a_{1},\ldots,$ $a_{m+n}$, $b_{1},\ldots,b_{m+n}$, $A,B,C,D$ with absolute value bounded by $T$ and integers $\gamma_{1},\gamma_{2}$ such that
$\left(C\tau_{1}+D\right)\tau_{2}=A\tau_{1}+B$
$\sum\limits_{p=1}^{m}(a_{p}+b_{p}\rho)z_{p}+(C\tau_{1}+D)\sum\limits_{q=1}^{n}(a_{m+q}+b_{m+q}\rho)w_{q}=\gamma_{1}+\gamma_{2}\tau_{1}$
And since $\lvert\tau_{1}\rvert,\lvert\tau_{2}\rvert,\lvert A\rvert,\lvert B\rvert,\lvert C\rvert,\lvert D\rvert,\lvert a_{1}\rvert,\ldots,\lvert a_{m+n}\rvert,\left|b_{1}\right|,\ldots,\left|b_{m+n}\right|\leq T$ and $z_{p}\in\mathcal{L}_{\tau_{1}},w_{q}\in\mathcal{L}_{\tau_{2}}$ we have that
$\left|\sum\limits_{p=1}^{m}(a_{p}+b_{p}\rho)z_{p}+(C\tau_{1}+D)\sum\limits_{q=1}^{n}(a_{m+q}+b_{m+q}\rho)w_{q}\right|$
$\leq\sum\limits_{p=1}^{m}\left(\lvert a_{p}\rvert+\left|b_{p}\right|\left|\rho\right|\right)\lvert z_{p}\rvert+\lvert C\tau_{1}+D\rvert\sum\limits_{q=1}^{n}\left(\lvert a_{m+q}\rvert+\left|b_{m+q}\right|\left|\rho\right|\right)\lvert w_{q}\rvert$
$\ll T\cdot\max\left\{1,\lvert\tau_{1}\rvert\right\}+(T\cdot\lvert\tau_{1}\rvert)\cdot T\cdot\max\left\{1,\lvert\tau_{2}\rvert\right\}\ll T^{4}$
and therefore we can assume that $\gamma_{1}$ and $\gamma_{2}$ have absolute value $\ll T^{4}$. This in turn implies that
$\left(a_{1},\ldots,a_{m+n},b_{1},\ldots,b_{m+n},A,B,C,D,\gamma_{1},\gamma_{2},\mathrm{Re}(\tau_{1}),\mathrm{Im}(\tau_{1}),\mathrm{Re}(z_{1}),\mathrm{Im}(z_{1}),\ldots,\right.$
$\left.\mathrm{Re}(z_{m}),\mathrm{Im}(z_{m}),\mathrm{Re}(\tau_{2}),\mathrm{Im}(\tau_{2}),\mathrm{Re}(w_{1}),\mathrm{Im}(w_{1}),\ldots,\mathrm{Re}(w_{n}),\mathrm{Im}(w_{n})\right)\in W^{\sim}(\mathbb{Q},\delta T^{4})$
for some positive constant $\delta$. Then, by Lemma 4.3, for every element of $\pi_{2}(W^{\sim}(\mathbb{Q},\delta T^{4}))$ there are at most $c^{\prime}$ different elements of $\mathcal{Z}(T)$. Finally, we conclude the proof using Lemma 4.2. ∎
5. Arithmetic bounds
Let $\mathcal{C}$ as in Theorem 1.2 and let $\mathcal{C}^{\prime}$ be the set of points $\mathbf{c}\in\mathcal{C}(\mathbb{C})$ such that there exists an isogeny $\phi_{\mathbf{c}}:E_{\mu(\mathbf{c})}\rightarrow E_{\lambda(\mathbf{c})}$ and $\mathbf{a},\mathbf{b}\in\mathbb{Z}^{m+n}$ with $(a_{1}+b_{1}\rho,\ldots,a_{m+n}+b_{m+n}\rho)\neq\mathbf{0}$ and
$(a_{1}+b_{1}\rho)P_{1}(\mathbf{c})+\ldots+(a_{m}+b_{m}\rho)P_{m}(\mathbf{c})+(a_{m+1}+b_{m+1}\rho)\phi_{\mathbf{c}}(Q_{1}(\mathbf{c}))+\ldots+(a_{m+n}+b_{m+n}\rho)\phi_{\mathbf{c}}(Q_{n}(\mathbf{c}))=O$
where $\rho$ is 0 in cases (i) and (ii), and a fixed generator for $\operatorname{End}(E_{\lambda_{0}})$ in case (iii). Moreover, we can also assume that $\phi_{\mathbf{c}}$ is a cyclic isogeny.
Since $\mathcal{C}$ is defined over $\overline{\mathbb{Q}}$, the curve $\widetilde{\mathcal{C}}=(J\circ\pi)(\mathcal{C})$ is also defined over $\overline{\mathbb{Q}}$ and thus, for every $\mathbf{c}\in\mathcal{C}^{\prime}$, $\left(J(\lambda(\mathbf{c})),J(\mu(\mathbf{c}))\right)\in\widetilde{\mathcal{C}}\cap\bigcup_{N\geq 1}Y_{0}(N)$. As all the modular curves $Y_{0}(N)$ are defined over $\mathbb{Q}$, all the points $\left(J(\lambda(\mathbf{c})),J(\mu(\mathbf{c}))\right)$ are algebraic, which implies that also $\lambda(\mathbf{c})$ and $\mu(\mathbf{c})$ are algebraic for every $\mathbf{c}\in\mathcal{C}^{\prime}$. From this, it follows that $\mathcal{C}^{\prime}$ is a subset of $\mathcal{C}(\overline{\mathbb{Q}})$.
Lemma 5.1.
Let $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ and define $D_{0}=\left[k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0})):k\right]$. Then, there exists a positive constant $\gamma_{1}$ (depending only on $\mathcal{C}$) such that
$h(\lambda(\mathbf{c}_{0})),h(\mu(\mathbf{c}_{0}))\leq\gamma_{1}D_{0}.$
Proof.
Fix $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ and call $N_{0}$ the smallest integer such that there exists an isogeny $\phi_{\mathbf{c}_{0}}:E_{\mu(\mathbf{c}_{0})}\rightarrow E_{\lambda(\mathbf{c}_{0})}$ of degree $N_{0}$ (note that this implies that $\phi_{\mathbf{c}_{0}}$ is a cyclic isogeny). Then $(J(\lambda(\mathbf{c}_{0})),J(\mu(\mathbf{c}_{0})))\in(\widetilde{\mathcal{C}}\cap Y_{0}(N))(\overline{\mathbb{Q}})$ and, since $\widetilde{\mathcal{C}}$ is asymmetric, we can apply [Hab10, Theorem 1.1] and standard properties of heights to get
$h(\lambda(\mathbf{c}_{0})),h(\mu(\mathbf{c}_{0}))\leq\gamma_{5}\log(1+N_{0}).$
Finally, using a result of Pellarin [Pel95] we get that $N_{0}\ll D_{0}^{4}(h(E_{\lambda(\mathbf{c}_{0})}))^{2}$, where
$h(E_{\lambda(\mathbf{c}_{0})})=\max(1,h(g_{2}(\lambda(\mathbf{c}_{0}))),h(g_{3}(\lambda(\mathbf{c}_{0}))))\ll h(\lambda(\mathbf{c}_{0}))$
since $g_{2}(\lambda(\mathbf{c}_{0})),g_{3}(\lambda(\mathbf{c}_{0}))$ are polynomials in $\lambda(\mathbf{c}_{0})$ (see equation (3.7) in [MZ10]). So,
$N_{0}\ll D_{0}^{4}(h(E_{\lambda(\mathbf{c}_{0})}))^{2}\ll D_{0}^{4}h(\lambda(\mathbf{c}_{0}))^{2}\ll D_{0}^{4}\log^{2}(1+N_{0})\ll D_{0}^{4}N_{0}^{\frac{1}{2}}$
which implies that $N_{0}\ll D_{0}^{8}$. Combining this with the previous estimates concludes the proof. ∎
Remark 5.2.
Note that this lemma above is the only part of the proof where we need to use the hypothesis that $\mathcal{C}$ is asymmetric, while all the other steps are true also for non-asymmetric curves. Thus, if one was able to prove this lemma for an arbitrary $\mathcal{C}$ or any of the Conjectures 21.20, 21.23 or 21.24 from [Pil22], then Theorem 1.2 would follow for any $\mathcal{C}$.
Lemma 5.3.
Let $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ and define $D_{0}=\left[k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0})):k\right]$. Then, there exist positive constants $\gamma_{2},\gamma_{3},\gamma_{4}$ (depending only on $\mathcal{C}$) such that
$\widehat{h}(P_{j}(\mathbf{c}_{0}))\leq\gamma_{2}D_{0}\quad\text{for every }j=1,\ldots,m$
$\widehat{h}(\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0})))\leq\gamma_{3}D_{0}^{9}\quad\text{for every }j=1,\ldots,n.$
Moreover, the $P_{j}(\mathbf{c}_{0})$ and the $\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0}))$ are defined over a field $K\supseteq k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0}))$ with
$\left[K:\mathbb{Q}\right]\leq\gamma_{4}D_{0}^{17}.$
Proof.
We use the same notation as in the previous proof.
Using work of Zimmer [Zim76] and the previous lemma, in case (i) we have
$\widehat{h}(P_{j}(\mathbf{c}_{0}))\leq h(P_{j}(\mathbf{c}_{0}))+\gamma_{6}\left(h(\lambda(\mathbf{c}_{0}))+1\right)\leq\gamma_{7}\left(h(\lambda(\mathbf{c}_{0}))+1\right)\leq\gamma_{2}D_{0}$
while in case (ii) and (iii) we get the same estimate by
$\widehat{h}(P_{j}(\mathbf{c}_{0}))\leq h(P_{j}(\mathbf{c}_{0}))+\gamma_{6}\left(h(\mu(\mathbf{c}_{0}))+1\right)\leq\gamma_{7}\left(h(\mu(\mathbf{c}_{0}))+1\right)\leq\gamma_{2}D_{0}.$
Similarly, $\widehat{h}(Q_{j}(\mathbf{c}_{0}))\leq\gamma_{8}D_{0}$. So, by Proposition 2.3, we get that
$\widehat{h}(\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0})))\leq\gamma_{8}N_{0}D_{0}\leq\gamma_{3}D_{0}^{9}$
Lemma 6.1 in [BC16] implies that the $P_{j}(\mathbf{c}_{0})$ and the $Q_{j}(\mathbf{c}_{0})$ are defined over a field $K_{1}$ of degree $\leq\gamma_{9}D_{0}$ over $\mathbb{Q}$. However, since $\phi_{\mathbf{c}_{0}}$ is cyclic of degree $N_{0}$, $\phi_{\mathbf{c}_{0}}$ is defined over a field $K_{2}$ of degree $\ll N_{0}^{2}$ over $k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0}))$, by Vélu’s formulas [Vél71]. Therefore, the points $\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0}))$ are defined over the compositum $K_{1}K_{2}$ which has degree $\ll D_{0}N_{0}^{2}\ll D_{0}^{17}$ over $\mathbb{Q}$. ∎
Next, we show that for any $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ we can choose “small” coefficients $a_{i}\in\mathbb{Z}$ for the linear relation.
Lemma 5.4.
For any $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$, there exist $\mathbf{a},\mathbf{b}\in\mathbb{Z}^{m+n}$ with $(a_{1},\ldots,a_{m+n},b_{1},\ldots,b_{m+n})\neq\mathbf{0}$ and
$(a_{1}+b_{1}\rho)P_{1}(\mathbf{c}_{0})+\ldots+(a_{m}+b_{m}\rho)P_{m}(\mathbf{c}_{0})+(a_{m+1}+b_{m+1}\rho)\phi_{\mathbf{c}_{0}}(Q_{1}(\mathbf{c}_{0}))+\ldots+(a_{m+n}+b_{m+n}\rho)\phi_{\mathbf{c}_{0}}(Q_{n}(\mathbf{c}_{0}))=O$
and such that
$\max\left\{\lvert a_{i}\rvert,\left|b_{i}\right|\right\}\leq\gamma_{10}D_{0}^{\eta_{1}}$
for some positive constants $\gamma_{10},\eta_{1}$ depending only on $\mathcal{C}$ and $m+n$.
Proof.
For cases (i) and (ii), we already saw that we can take $\rho=0$ and therefore we can choose $\mathbf{b}=\mathbf{0}$. So the result is a direct consequence of Lemma 6.1 of [BC16] (which is in turn based on a result by Masser [Mas88]), applied to the points $P_{i}(\mathbf{c}_{0})$ and $\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0}))$, and the height bounds from the previous lemma.
In case (iii), we use again the above-mentioned lemma by Barroero and Capuano, this time with the points $P_{i}(\mathbf{c}_{0})$, $\rho P_{i}(\mathbf{c}_{0})$, $\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0}))$ and $\rho\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0}))$, recalling that by (2.3), we have that
$\widehat{h}(\rho P_{i}(\mathbf{c}_{0}))\ll D_{0}^{6}\cdot\widehat{h}(P_{i}(\mathbf{c}_{0}))\ll D_{0}^{7}$
$\widehat{h}(\rho\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0})))\ll D_{0}^{6}\cdot\widehat{h}(\phi_{\mathbf{c}_{0}}(Q_{j}(\mathbf{c}_{0})))\ll D_{0}^{15}.$
For the next lemma, let $\tau_{1}(\mathbf{c})=\tau_{1}(\varphi^{-1}(\mathbf{c}))\in\mathcal{B}$ for every $\mathbf{c}\in\mathcal{C}(\mathbb{C})$ and similarly for $\tau_{2}(\mathbf{c})$, where $\varphi$ was defined in section 2.2.
Lemma 5.5.
There exists a positive constant $\gamma_{11}$, depending only on $\mathcal{C}$, such that for every $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ we have
$\lvert\tau_{1}(\mathbf{c}_{0})\rvert,\lvert\tau_{2}(\mathbf{c}_{0})\rvert\leq\gamma_{11}D_{0}^{2}$
Proof.
Since $\mathcal{B}$ was defined as the union of six fundamental domains for the action of $\mathrm{SL}_{2}(\mathbb{Z})$ on $\mathbb{H}$, we only need to prove the result for $\tau_{1,2}(\mathbf{c}_{0})$ in the usual fundamental domain $\mathcal{D}$.
For $\tau\in\mathcal{D}$, Lemma 1 in [BMZ13] implies that $e^{2\pi\mathrm{Im}(\tau)}\leq 2079+\left|j(\tau)\right|$. Hence, if $\left|j(\tau)\right|\leq 2$, then $\mathrm{Im}(\tau)\leq\frac{1}{2\pi}\log(2081)=\gamma_{12}$. Equivalently, for every $\tau\in\mathcal{D}$ such that $\mathrm{Im}(\tau)>\gamma_{12}$, then $\left|j(\tau)\right|>2$.
If $\mathrm{Im}(\tau)\leq\gamma_{12}$, then $\left|\tau\right|\leq\sqrt{\frac{1}{4}+\gamma_{12}^{2}}$, so we may assume $\gamma_{11}\geq\sqrt{\frac{1}{4}+\gamma_{12}^{2}}$.
Suppose now that $\mathrm{Im}(\tau)>\gamma_{12}$, we then get that
$\mathrm{Im}(\tau)\leq\dfrac{1}{2\pi}\log\left(2079+\left|j(\tau)\right|\right)\leq\dfrac{\log(2081)}{2\pi\log(2)}\log\left|j(\tau)\right|$
Moreover, we have that $j(\tau_{1}(\mathbf{c}_{0}))=J(L(\tau_{1}(\mathbf{c}_{0})))=J(\lambda(\mathbf{c}_{0}))$, where $J(\lambda)=2^{8}\frac{(\lambda^{2}-\lambda+1)^{3}}{\lambda^{2}(\lambda-1)^{2}}$ and $L$ was defined in (2.1). Since $\lambda(\mathbf{c}_{0})\in\overline{\mathbb{Q}}\setminus\left\{0,1\right\}$, this implies that $j(\tau_{1}(\mathbf{c}_{0}))\in\overline{\mathbb{Q}}$.
Then, using the inequality $\log\left|\alpha\right|\leq\left[\mathbb{Q}(\alpha):\mathbb{Q}\right]h(\alpha)$ for every non-zero $\alpha\in\overline{\mathbb{Q}}$, we get
$\log\left|j(\tau_{1}(\mathbf{c}_{0})\right|$ $\leq\left[\mathbb{Q}(j(\tau_{1}(\mathbf{c}_{0}))):\mathbb{Q}\right]h(j(\tau_{1}(\mathbf{c}_{0})))=\left[\mathbb{Q}(J(\lambda(\mathbf{c}_{0}))):\mathbb{Q}\right]h(J(\lambda(\mathbf{c}_{0})))$
$\leq\gamma_{13}\left[\mathbb{Q}(\lambda(\mathbf{c}_{0})):\mathbb{Q}\right](h(\lambda(\mathbf{c}_{0}))+1)\leq\gamma_{14}D_{0}^{2}$
by Lemma 5.1. Combining this with the previous bound gives $\mathrm{Im}(\tau_{1}(\mathbf{c}_{0}))\leq\gamma_{15}D_{0}^{2}$ and we easily get the estimate in the statement. ∎
6. Proof of Theorem 1.2
We want to show that the set $\mathcal{C}^{\prime}$ is finite. By Northcott’s theorem and Lemma 5.1, we only need to bound the degree $D_{0}$ of $\lambda(\mathbf{c})$ and $\mu(\mathbf{c})$ over $k$.
Let $\mathbf{c}_{0}\in\mathcal{C}^{\prime}$ and $\sigma\in\operatorname{Gal}(\overline{k}/k)$. Notice that $\sigma(\mathbf{c}_{0})\in\mathcal{C}^{\prime}$, since
$j\left(E_{\lambda(\sigma(\mathbf{c}_{0}))}\right)=j\left(E_{\sigma(\lambda(\mathbf{c}_{0}))}\right)=J\left(\sigma(\lambda(\mathbf{c}_{0}))\right)=\sigma\left(J\left(\lambda(\mathbf{c}_{0})\right)\right)=\sigma\left(j\left(E_{\lambda(\mathbf{c}_{0})}\right)\right)$
implies that
$\Phi_{N_{0}}\left(j\left(E_{\lambda(\sigma(\mathbf{c}_{0}))}\right),j\left(E_{\mu(\sigma(\mathbf{c}_{0}))}\right)\right)=\sigma\left(\Phi_{N_{0}}\left(j\left(E_{\lambda(\mathbf{c}_{0})}\right),j\left(E_{\mu(\mathbf{c}_{0})}\right)\right)\right)=0$
and also
$P_{i}(\sigma(\mathbf{c}_{0}))=\sigma\left(P_{i}(\mathbf{c}_{0})\right)$
$\phi_{\sigma(\mathbf{c}_{0})}\left(Q_{i}(\sigma(\mathbf{c}_{0}))\right)=\phi_{\sigma(\mathbf{c}_{0})}\left(\sigma\left(Q_{i}(\mathbf{c}_{0})\right)\right)=\sigma\left(\phi_{\mathbf{c}_{0}}\left(Q_{i}(\mathbf{c}_{0})\right)\right).$
Moreover, in case (iii), we can assume without loss of generality that the generator $\rho$ of $\operatorname{End}(E_{\lambda_{0}})$ is in $k$, so that $\sigma(\rho)=\rho$.
So, in all cases:
$(a_{1}+b_{1}\rho)$ $P_{1}(\sigma(\mathbf{c}_{0}))+\ldots+(a_{m}+b_{m}\rho)P_{m}(\sigma(\mathbf{c}_{0}))+$
$+(a_{m+1}+b_{m+1}\rho)\phi_{\sigma(\mathbf{c}_{0})}(Q_{1}(\sigma(\mathbf{c}_{0})))+\ldots+(a_{m+n}+b_{m+n}\rho)\phi_{\sigma(\mathbf{c}_{0})}(Q_{n}(\sigma(\mathbf{c}_{0})))=$
$=\sigma\Big{(}(a_{1}+b_{1}$ $\rho)P_{1}(\mathbf{c}_{0})+\ldots+(a_{m}+b_{m}\rho)P_{m}(\mathbf{c}_{0})+$
$+(a_{m+1}+b_{m+1}\rho)\phi_{\mathbf{c}_{0}}(Q_{1}(\mathbf{c}_{0}))+\ldots+(a_{m+n}+b_{m+n}\rho)\phi_{\mathbf{c}_{0}}(Q_{n}(\mathbf{c}_{0}))\Big{)}=O$
on $E_{\lambda(\sigma(\mathbf{c}_{0}))}$, since the $a_{i}$ and $b_{i}$ are integers.
Now, consider the point $\varphi^{-1}\left(\sigma(\mathbf{c}_{0})\right)\cap\mathcal{F}_{\mathcal{B}}\in\mathcal{Z}$ with coordinates $\left(\tau_{1}^{\sigma},z_{1}^{\sigma},\ldots,z_{m}^{\sigma},\tau_{2}^{\sigma},w_{1}^{\sigma},\ldots,w_{n}^{\sigma}\right)$ (here the superscript $\sigma$ does not denote a Galois conjugate). By the previous equation and lemmas 5.4 and 5.5 we have relations
$(a_{1}+b_{1}\rho)z_{1}^{\sigma}+\ldots+(a_{m}+b_{m}\rho)z_{m}^{\sigma}+(C\tau_{1}^{\sigma}+D)\left((a_{m+1}+b_{m+1}\rho)w_{1}^{\sigma}+\ldots+(a_{m+n}+b_{m+n}\rho)w_{n}^{\sigma}\right)=\mathbb{Z}+\tau_{1}^{\sigma}\mathbb{Z}$
$\tau_{2}^{\sigma}=\dfrac{A\tau_{1}^{\sigma}+B}{C\tau_{1}^{\sigma}+D}$
with
$\max\left\{\left|a_{i}\right|,\left|b_{i}\right|\right\}\leq\gamma_{10}D_{0}^{\eta_{1}}\qquad\left|\tau_{1}^{\sigma}\right|,\left|\tau_{2}^{\sigma}\right|\leq\gamma_{11}D_{0}^{2}$
and $\left|A\right|,\left|B\right|,\left|C\right|,\left|D\right|\leq\gamma_{16}N_{0}^{10}\leq\gamma_{17}D_{0}^{80}$ by Lemma 2.1 and Lemma 5.1.
So,
$\varphi^{-1}\left(\sigma(\mathbf{c}_{0})\right)\cap\mathcal{F}_{\mathcal{B}}\in\mathcal{Z}\left(\gamma D_{0}^{\eta}\right)$
where $\gamma=\max\left\{\gamma_{10},\gamma_{11},\gamma_{17}\right\}$ and $\eta=\max\left\{\eta_{1},80\right\}$.
There are at least $\left[k(\mathbf{c}_{0}):k\right]\geq\left[k(\lambda(\mathbf{c}_{0}),\mu(\mathbf{c}_{0})):k\right]=D_{0}$ different $\left(\tau_{1}^{\sigma},z_{1}^{\sigma},\ldots,z_{m}^{\sigma},\tau_{2}^{\sigma},w_{1}^{\sigma},\ldots,w_{n}^{\sigma}\right)$ in $\mathcal{Z}\left(\gamma D_{0}^{\eta}\right)$. However, applying Proposition 4.1 with $\varepsilon=\frac{1}{2\eta}$ gives a contradiction if $D_{0}$ is large enough. This proves that $D_{0}$ is bounded and, consequently, Theorem 1.2.
Acknowledgements
We would like to thank Fabrizio Barroero and Laura Capuano for many useful discussions and comments, and Gabriel Dill for his help, for having pointed out the alternative proof of cases (ii) and (iii) and for his hospitality in Bonn.
The author was supported by the PRIN 2022 project 2022HPSNCR: Semiabelian varieties, Galois representations and related Diophantine problems and the National Group for Algebraic and Geometric Structures, and their Applications (GNSAGA INdAM).